# An Overview of Some Single Machine Scheduling Problems: Polynomial Algorithms, Complexity and Approximability

### v1

@@ 1 Introduction
In our everyday life we often wish to optimize out time, to do our duties in a due time or to arrange as many things as possible in a limited time. Sometimes optimization of our time is a purely personal matter, but often it depends on the factors and the environment around us. We wish to arrange them in a way which would be beneficial for us. As an example, we give a problem which many of us encounter daily. A family lives in an apartment with a single bathroom. Each member of the family daily gets up at a fixed time, needs prescribed time units in the bathroom for the prescribed time units. The bathroom is our resource (machine) and can be used by at most one person at a time. Since each member of the family has to leave home at a prescribed time (in order the reach the job/school on time), we face an optimization problem, how to arrange the activities of each member of the family on the scared resource to minimize the overall time needed for all members of the family to complete all activities. These types of problems are dealt with in discrete optimization in scheduling theory.
In general terms, a scheduling problem deals with a finite set of requests called jobs or tasks which have to be scheduled (performed) on a resource called machine or processor. In our family example, to go to the bathroom is an example of a request; the bathroom is an example of a resource. A job in a factory or a program in a computer system or a lesson at school are other examples of requests. A machine in a factory or a processor in a computer system or a teacher in a school are other examples of resources. Each job has its processing time, i.e., it needs a prescribed time on the machine, and the machine cannot handle more than one request at a time (for example, two teacher cannot give different lessons in a classroom simultaneously). Besides, the time when the resource is available is usually limited. We need to arrange an order in which the jobs are handled by the machine to minimize or maximize some important, often time, criterion or objective functions. A common objective function is the maximal job completion time or the makespan usually denoted by $C_{\max}$.
@@
We shall denote the given set of $n$ jobs by $\{1,2,\dots,n\}$. The processing time of job $j$ is denoted by $p_{j}$. Job $j$ may have some additional characteristics. The release time (or head) $r_{j}$ is the time moment when job $j$ becomes available (it cannot start before this time). The delivery time (or tail) $q_{j}$ is the time which takes $j$ after it is completely processed (completed) on the machine (for example, a job should be delivered to a remote customer). The job delivery takes no machine time (a job is delivered by an independent agent to a customer), though note that $q_{j}$ contributes to the completion time of job $j$, hence the makespan crucially depends on the job delivery times. The due date $d_{j}$ is the desirable completion time of job $j$. Both above parameters specify the urgency of job $j$: the more is $q_{j}$ more urgent is job $j$; likewise, the less is $d_{j}$, the more urgent is job $j$.
@@
All above parameters are normally non-negative integral numbers. In some cases, job $j$ is not allowed to be completed after its due date. Then the term deadline instead of “due date” is commonly used. Thus, if deadlines are given, our objective is to find a schedule in which all jobs are completed before or at their deadlines. With due dates, our objective will be to minimize or maximize some due date oriented objective function. Let the lateness of a job in a schedule be the difference between its completion time and its due date. One of the common objective functions is the maximal job lateness $L_{\max}$ (which we have to minimize). Another common objective function is the number of jobs completed at or before their deadlines $\sum U_{j}$ which we have to maximize ($U_{j}$ is a 0-1 variable indicating whether $j$ is completed or not in the schedule). In some real-time overloaded systems, the total workload often exceeds the capacity of the processor. Job due dates are critical, in the sense that the jobs that are not completed by their due dates bring little benefit and might be completed arbitrarily late. In such systems, a reasonable goal is to maximize the number of jobs completed by their due dates, or more generally, to maximize their value or profit. The profit of job $j$ is called weight and is denoted by $w_{j}$.
Job preemptions might be allowed or not. In a preemptive scheduling problem, a job can be interrupted on the machine and later resumed on the machine. In a non-preemptive scheduling problem, a job $J_{j}$, once scheduled on the machine, should be processed without any interruption until its full completion on the machine. The machine is called idle at some moment $t$ if no job is processed by the machine at that moment, otherwise it is said to be busy.
Certain restrictions on how the jobs can be scheduled on the machine define the whole set of all feasible schedules. One of the principle restrictions are resource (machine) restrictions: the machine can handle no more than one job at a time. Likewise, we can have precedence relations between the jobs, i.e., the set of jobs can be partially ordered (some jobs cannot be started before the others are not completed). Both type of restrictions imply that the jobs have to be processed in a sequential manner. The precedence restrictions are traditionally represented by directed graphs, the so-called precedence (task) graphs. In such graphs nodes represent jobs, and there is an arc $(i,j)$ if and only if job $i$ immediately precedes job $j$.
@@
There exist a huge number of papers dealing with single machine problems. In particular, many heuristics and metaheuristic as well as enumerative algorithms have been suggested and tested for special single machine problems. Regarding surveys, the authors often restrict to special subjects and/or consider more general problem settings. As an example, there exist three very detailed surveys on scheduling problems with setup times or costs, see [4, 5, 6]. For problems with so-called $p$-batching (i.e., several jobs can be processed in parallel in a batch), a survey has been given by Fowler and Münch [17]. The case of equal processing times has been surveyed by Kravchenko and Werner [31].
@@ 2 Some Basics
A gap in a schedule is a maximal consecutive time interval in which no job is processed by the machine. It can be assumed that there occurs a 0-length gap $(c_{j},t_{i})$ if job $i$ starts at its release time at the the completion time of job $j$.
A block in a schedule is its consecutive part which is preceded and succeeded by a gap.
@@
If both, job release and due dates are given, then it makes sense to minimize the maximum job lateness (note that due dates are irrelevant for the makespan minimization). First, it is useful to note that, if we allow preemptions, the following simple on-line strategy is optimal. Starting from time 0, at any time $t$, among all available jobs, schedule next the one with the smallest due date. If a job with a due date smaller than the currently running job becomes available, then interrupt the latter job and schedule a newly released job with the smallest due date. The resulting schedule will have no machine idle time unless no job is available.
Assume that job $i$ is processed at time $r_{j}$ in the schedule $S$, where $d_{i}>d_{j}$. Without allowing the interruption of $i$, this job delays the starting of job $j$; note that this delay will be less than $p_{i}$ since upon the completion of job $i$, job $j$ can immediately be scheduled. The delay of job $j$ may result in a non-optimal schedule. Indeed, it may turn out that job $j$ has a “small” due-date so that it had to be scheduled without or with a less delay in an optimal schedule (the delay of a job $j$ in the schedule $S$ is $t_{j}^{S}-r_{j}$).
Now we describe a basic popular greedy algorithm that has been widely used. This earliest due-date (ED) heuristic) was suggested by Jackson [30] and later developed by Schrage [42] long time ago. The first version is just the above described greedy algorithm with the preprocessing stage. The second version was adopted for jobs having both, release and due times. We describe this heuristic below
Let $t$ be the maximum between the minimal release time ofa yet unscheduled job and the time when the machine completes the latest scheduled job. Let this time be 0 if no job is yet scheduled. Initially, we have no scheduled jobs. Iteratively, the ED-heuristic, among all available jobs at time $t$, schedules a job with the smallest due date breaking ties by selecting a longest job. Observe that this heuristic is on-line and non-preemptive.
Let $\sigma$ be the initial ED-schedule, the one generated by the ED-heuristic for the originally given problem instance. Since no urgent job can be delayed by more than the processing time of another job, it is easy to see that
@@
If the ED-schedule $S$ is not optimal, the overflow job $o(S)$ is to be restarted earlier. By modifying artificially the initial problem parameters and applying the ED-heuristic to this modified instance, an alternative ED-schedule in which job $o(S)$ is restarted earlier can be created, as we describe now. In the schedule $S$, we look for the overflow job $o(S)$, the kernel $K(S)$ and the set of all emerging jobs. In an attempt to reduce $L_{\max}(S)$ (restart job $o(S)$ earlier), we activate an emerging job $l$ for the kernel $K(S)$, that is, we reschedule job $l$ after the kernel $K(S)$. Technically, we accomplish this in two steps: First, we increase artificially the release time of job $l$, assigning to it a magnitude, no less than the largest job release time in the kernel $K(S)$. Then we apply the ED-heuristic to the modified problem instance. Since now the release time of job $l$ is no less than that of any job of the kernel $K(S)$, and $d_{l}$ is larger than the due date of any job from the kernel $K(S)$, the ED-heuristic will give the priority to the jobs of the kernel $K(S)$ and will reschedule job $l$ after all former jobs. The ED-schedule $\{S\}_{l}$, obtained in this way, is called a (weakly) complementary to $S$ schedule.
Note that in the schedule $S$, some emerging jobs might have been originally included after the kernel $K(S)$. Then at least one such job will be included before the jobs of the kernel $K(S)$ in the schedule $\{S\}_{l}$ by the ED-heuristic (since no job of the kernel $K(S)$ is released at time $s_{l}(S)$). As a result, the kernel $K(S)$ may not be restarted at its earliest starting time. To guarantee the restarting of the kernel $K(S)$ at its earliest starting time, the above emerging jobs might be forced to remain after the kernel $K(S)$. We can clearly achieve this by increasing the release times of these jobs, similarly as we did it for job $l$. The ED-schedule, obtained by the ED-heuristic for the modified in this way problem instance, is called a (strongly) complementary to $S$ schedule and is denoted by $S_{l}$. A complementary schedule can be abbreviated as a C-schedule.
@@
Let $l$ be the emerging job scheduled in $i$th position in a strongly C-schedule $S$ and let $j$ be the job, immediately succeeding $l$ in $S$. Then there will arise a gap somewhere between the $i$th scheduled job and the kernel $K(S)$ in the schedule $S_{l}$ if $d_{j}<d_{l}$. In particular, if $l$ is the live emerging job in the schedule $S$, there will arise a gap immediately before the earliest scheduled job of the kernel $K(S)$ in the complementary schedule $S_{l}$, and this kernel will start at its earliest starting time.
@@ 3.1 Scheduling a Single Machine with Release Times and Due Dates to Minimize Maximum Job Lateness
The general problem with release tames and due dates to minimize maximum lateness, abbreviated as $1|r_{j},d_{j}|L_{\max}$, is strongly NP-hard [23]. If we replace the due dates by tails with the objective to minimize maximum lateness, we get an equivalent problem $1|r_{j},q_{j}|C_{\max}$. In this transformation, every due date $d_{j}$ is replaced by the tail $q_{j}=A-d_{j}$, where $A$ is a sufficiently large constant (no less than the maximum job due date). So, the smaller is the due date, the larger is the tail. In other words, the urgency of every job is conserved with this transformation. It is straightforward to see that the two versions are equivalent. The full completion time of job $j$ in $S$ is given by ${\cal C}_{j}(S)=c_{j}^{S}+q_{j}$. The objective in that setting is to find an optimal makespan schedule, i.e., a feasible schedule $S$ with the minimal value of the maximal full job completion time.
An efficient implicit enumerative algorithms for the problems $1|r_{j},d_{j}|L_{\max}$ and $1|r_{j},q_{j}|C_{\max}$ was suggested in McMahon and Florian [38] and later, based on similar ideas, in Carlier [13]. A number of other special cases are solvable in polynomial time. If all jobs are of unit length, the problem is again easy to solve, see Horn [29]: indeed, while applying the ED-heuristic, no new (earlier unreleased) job can be released within the execution interval of a currently running job since job release times are integer and every job takes just a unit time of processing. As a result, no job may cause a forced delay of any more urgent job.
@@
With arbitrary release times and due dates (or tails), if all processing times are equal (but not necessarily of unit length), it is still possible to find an optimal schedule in polynomial time Garey et al [24]. Even with two allowable job processing times $p$ and $2p$, for an arbitrary integer $p$, the problem $1|p_{j}\in\{p,2p\},r_{j},q_{j}|C_{\max}$ remains polynomially solvable [45]. In fact, in the algorithm from the latter reference the processing times of only some jobs (which become known during the execution of the algorithm) are restricted to $p$ or $2p$, whereas the rest of the jobs may have arbitrary processing times. The algorithm is obtained as a consequence of two enumerative algorithms for the problem $1|r_{j},q_{j}|C_{\max}$. The first algorithm is exponential, whereas the second one, which is a restriction of the first one, runs in $O(n^{2}\log n)$ time if certain conditions during its execution are satisfied. Otherwise, another auxiliary procedure is to be applied to guarantee the optimality. An $O(n^{2}\log n\log P)$ implementation of that procedure for the problem $1|p_{j}\in\{P,2P\},r_{j},q_{j}|C_{\max}$ yields an algorithm with the same time complexity for this setting.
@@
$r_{i}+p_{i}\geq r_{j}+p_{j}$
$d_{i}\geq d_{j}$
holds, then the resultant problem $1|r_{j},embedded|L_{\max}$ is solvable in $O(n\log n)$ time, see Vakhania [49].
@@
With arbitrary job processing times, the problem remains (weakly) $NP$-hard even with only two allowable job release times and due dates $1|r_{j}\in\{r_{1},r_{2}\},d_{j}\in\{d_{1},d_{2}\}|L_{\max}$ [14]. This is a minimal $NP$-hard special case of the problem $1|r_{j},d_{j}|L_{\max}$.
@@
Finally, a special kind of parameterized analysis called the Variable Parameter (VP) analysis was recently suggested in [51]. Based on this approach, an implicit enumeration algorithm and a polynomial time approximation scheme for the general setting $1|r_{j},q_{j}|C_{\max}$ were proposed. The variable parameter is defined as the number of emerging jobs. First, a partial solution in time $O(n\log n)$ without these jobs is constructed. Then this partial solution is augmented to a complete optimal solution in time, which is exponential solely in the number of the variable parameter (the number of emerging jobs). The author also gives alternative time complexity expression for both, the implicit enumeration and the polynomial time approximation scheme, in which the exponential dependence is solely on the maximum processing time and the maximum delivery time of an emerging job. If one applies the fixed parameter analysis to these estimations, one gets a polynomial-time dependence. A very recent extensive experimental study showed that, in practice, the variable parameter is much less than $n$. Moreover, the ratio of the variable parameter to $n$ asymptotically converges to 0 (see [41]).

### v2

@@ 1 Introduction
Scheduling problems arise in many real-life applications including our everyday life. We commonly desire to optimize our time arranging different activities, where the time is limited and also the resources that we use are scarce. We wish to optimize some of our personal criteria but we also depend on the availability of the latter resources. As a simple but very common daily example, consider an apartment with a single bathroom, where a small family lives. The family members gets up at some fixed time, and all of them need to spend some time in the bathroom, which is a scarce resource (or machine). Typically such a (renewable) resource may serve a single object or job, a family member in our example. Every member of the family needs to leave the apartment at some fixed time to reach his/her job or school without delay. This time can be different for different members of the family. Here arises a small optimization problem that asks to schedule the time interval for each member of the family so as to meet the deadline of each member of the family.
Generally, in a scheduling problem we are given jobs or tasks which are some requests that need to be performed by some resource, a machine or a processor. In our small example, bathroom requirements for each member of the family are jobs and the bathroom is a machine. We can give a vast amount of examples of jobs and machines in our practical life; e.g., a job in a factory or a software application in a computer or on a smart phone, or a lesson at school. Here machines, computers, smart phones, classrooms and teachers are resource examples of. Every jobs requires some time units on a resource, whereas a resource is able to process at most one job at a time. Indeed, one classroom cannot be used for two different classes simultaneously or a a teacher cannot give two or more classes at the same time. There is a restriction on the whole time window during which a factory or a teacher can work. It is a scheduling task to arrange the order for processing the jobs. Often there is some minimization or maximization criterion called an objective function. For example, a very common criterion is to minimize the maximum job or machine completion time or the makespan which is commonly denoted by $C_{\max}$.
@@
We shall denote the given set of $n$ jobs by $\{1,2,\dots,n\}$. The processing time of job $j$ is denoted by $p_{j}$. Job $j$ may have some additional characteristics. For example, the time moment from when job $j$ is available to be processed is called its release or readiness time and is commonly denoted by $r_{j}$. In many applications, a job, once completed by the machine, needs to be delivered to the customer by an independent transportation unit, which may not be a scarce resource in a given model. Such a delivery time (or tail) of job $j$ is commonly denoted by $q_{j}$. The job delivery takes no machine time (a job is delivered by an independent agent to a customer), though $q_{j}$ contributes to the completion time of job $j$. Hence, the makespan crucially depends on the job delivery times. The due date $d_{j}$ is the desirable completion time of job $j$. Both above parameters specify the urgency of job $j$: the more is $q_{j}$, the more urgent is job $j$; likewise, the less is $d_{j}$, the more urgent is job $j$.
@@
All above parameters are normally non-negative integral numbers. In some cases, job $j$ is not allowed to be completed after its due date. Then the term deadline instead of “due date” is commonly used. Thus, if deadlines are given, our objective is to find a schedule in which all jobs are completed before or at their deadlines. With due dates, our objective will be to minimize or maximize some due date oriented objective function. Let the lateness of a job in a schedule be the difference between its completion time and its due date. One of the common objective functions is the maximal job lateness $L_{\max}$ (which we have to minimize). Another common objective function is the number of jobs completed at or before their due dates $\sum U_{j}$ which we have to maximize ($U_{j}$ is a 0-1 variable indicating whether job $j$ is completed on time or not in the schedule). We may allow to omit some jobs in the case when the total workload is more than the machine capacity. In many applications, job due dates are important since, if a job is not completed by its due date then its benefit or functionality becomes neglectable. Hence, we may wish to maximize the throughput, i.e.,the number of jobs completed on time by their due dates. We may also consider more general criteria to maximize some profits. The profit of job $j$ can be expressed by its weight which is commonly denoted by $w_{j}$. Job preemptions might be allowed or not. In a preemptive scheduling problem, a job can be interrupted on the machine and later resumed on the machine. In a non-preemptive scheduling problem, a job $J_{j}$, once scheduled on the machine, should be processed without any interruption until its full completion on the machine. The machine is called idle at some moment $t$ if no job is processed by the machine at that moment, otherwise it is said to be busy.
Problem restrictions on the way in which the jobs are to be performed by machines define the set of feasible solutions. For example, a machine can process at most one job simultaneously; the precedence relations defined on the set of jobs, restrict the relative order in which the jobs can be scheduled. Scheduling problems are sometimes referred to as sequencing problems since the jobs need to be scheduled in a sequential manner (as no job overlapping is permitted). The precedence relations between the jobs can be represented by directed graphs which are often referred to as precedence (task) graphs. In such graphs, nodes represent jobs, and there is an arc $(i,j)$ if and only if job $i$ immediately precedes job $j$.
@@
There exist a huge number of papers dealing with single machine problems. In particular, many heuristics and metaheuristic as well as enumerative algorithms have been suggested and tested for special single machine problems. Regarding surveys, the authors often restrict to special subjects and/or consider more general problem settings. As an example, there exist three very detailed surveys on scheduling problems with setup times or costs, see [4, 5, 6]. For problems with so-called $p$-batching (i.e., several jobs can be processed in parallel in a batch), a survey has been given by Fowler and Mönch [18]. The case of equal processing times has been surveyed by Kravchenko and Werner [32].
@@ 2 Some Basics
A gap on a machine is a longest time interval during which this machine is idle, i.e., it processes no job. It can be assumed that there occurs a 0-length gap $(c_{j},t_{i})$ if job $i$ starts at its release time at the the completion time of job $j$.
A block is a consecutive part in a schedule which is preceded and succeeded by a gap.
@@
If both, job release and due dates are given, then it makes sense to minimize the maximum job lateness (note that due dates are irrelevant for the makespan minimization). Suppose preemptions are allowed. Then we have a simple method to solve the problem optimally even on-line. We usually start at the earliest job release time and set initially the current time $t$ to that magnitude. Iteratively, among the available jobs, we schedule the next a most urgent job, one with the smallest due date. However, if during the execution of that job, another more urgent job gets released, the former job is interrupted at the release time of the latter job, which gets included at its release time. In this way, there will occur no forced delay for any urgent job by a non-urgent one, unlike the case without preemptions (in a given schedule $S$, the delay of job $j$ is $t_{j}^{S}-r_{j}$). The so constructed schedule is compact also in the sense that it minimizes the makespan. In particular, it has no avoidable machine idle time.
Assume that job $i$ is processed at time $r_{j}$ in the schedule $S$, where $d_{i}>d_{j}$. Without allowing the interruption of $i$, this job delays the starting of job $j$; note that this delay will be less than $p_{i}$ since upon the completion of job $i$, job $j$ can immediately be scheduled. The delay of job $j$ may result in a non-optimal schedule. Indeed, this will be the case when job $j$ is too urgent with a “small” due-date. In an optimal schedule, it should be included with no or with a smaller delay.
Now we describe a basic popular greedy non-preemptive algorithm to which we will refer as earliest due-date (ED) heuristic). It that has been widely used for solution of different single machine and other more complex scheduling problems. As the above described preemptive one, it is on-line. The heuristic was suggested by Jackson [31] and later developed by Schrage [45]. The first version is just the above described greedy algorithm with the pre-processing stage. The second version was adopted for jobs having both, release and due times. We describe this heuristic now. We let the current scheduling time $t$ be the minimal release time of an unscheduled job or the time when the machine completes the last job, whichever magnitude is larger. If no job is yet scheduled, we let $t:=0$. Iteratively, among the jobs already released by time $t$, the heuristic schedules one with the smallest due date. It breaks ties selecting a longest job.
Let $\sigma$ be the initial ED-schedule generated by the ED-heuristic for the originally given problem instance. Since no urgent job can be delayed by more than the processing time of another job, it is easy to see that
@@
If the ED-schedule $S$ is not optimal, the overflow job $o(S)$ is to be restarted earlier. By modifying artificially the initial problem parameters and applying the ED-heuristic to this modified instance, an alternative ED-schedule in which job $o(S)$ is restarted earlier can be created, as we describe now. In the schedule $S$, we look for the overflow job $o(S)$, the kernel $K(S)$ and the set of all emerging jobs. In an attempt to reduce $L_{\max}(S)$ (restart job $o(S)$ earlier), we activate an emerging job $l$ for the kernel $K(S)$, that is, we reschedule job $l$ after the kernel $K(S)$. The activation is carried out in two steps. At step 1 the release time of job $l$ is increased artificially to it a magnitude, no-less than the largest job release time in the kernel $K(S)$. At step, ED-heuristic is applied to the modified in this way problem instance. Since now the release time of job $l$ is no less than that of any job of the kernel $K(S)$, and $d_{l}$ is larger than the due date of any job from the kernel $K(S)$, the ED-heuristic will give the priority to the jobs of the kernel $K(S)$ and will reschedule job $l$ after all former jobs. The ED-schedule $\{S\}_{l}$, obtained in this way, is called a (weakly) complementary to $S$ schedule.
Note that in the schedule $S$, some emerging jobs might have been originally included after the kernel $K(S)$. Then at least one such job will be included before the jobs of the kernel $K(S)$ in the schedule $\{S\}_{l}$ by the ED-heuristic (since no job of the kernel $K(S)$ is released at time $s_{l}(S)$). As a result, the kernel $K(S)$ may not be restarted at its earliest starting time. To guarantee the restarting of the kernel $K(S)$ at its earliest starting time, the above emerging jobs might be forced to remain after the kernel $K(S)$. To reach this goal, we increase the release times of these jobs, similarly as we did it for job $l$. The ED-schedule, obtained by the ED-heuristic for the modified in this way problem instance, is called a (strongly) complementary to $S$ schedule and is denoted by $S_{l}$. A complementary schedule can be abbreviated as a C-schedule.
@@
Let $S$ be a strongly C-schedule and let $l$ be the emerging job scheduled in $i$th position in in that schedule. Let, further $j$ be the job, immediately succeeding job $l$ in schedule $S$. Then if $d_{j}<d_{l}$, there will occur a gap in between the $i$th scheduled job and the kernel $K(S)$ in the schedule $S_{l}$ Furthermore, there will arise a gap immediately before the earliest scheduled job of the kernel $K(S)$ in the complementary schedule $S_{l}$, where $l$ is the live emerging job in the schedule $S$. As a result, the kernel $K(S)$ will start at the minimum release time of a job from that kernel.
@@ 3.1 Scheduling a Single Machine with Release Times and Due Dates to Minimize Maximum Job Lateness
The general problem with release tames and due dates to minimize maximum lateness, abbreviated as $1|r_{j}|L_{\max}$, is strongly NP-hard [24]. If we replace the due dates by tails with the objective to minimize maximum lateness, we get an equivalent problem $1|r_{j},q_{j}|C_{\max}$. In this transformation, every due date $d_{j}$ is replaced by the tail $q_{j}=A-d_{j}$, where $A$ is a sufficiently large constant (no less than the maximum job due date). So, the smaller is the due date, the larger is the tail. In other words, the urgency of every job is conserved with this transformation. It is straightforward to see that the two versions are equivalent. The full completion time of job $j$ in $S$ is given by ${\cal C}_{j}(S)=c_{j}^{S}+q_{j}$. The objective in that setting is to find an optimal makespan schedule, i.e., a feasible schedule $S$ with the minimal value of the maximal full job completion time.
An efficient implicit enumerative algorithms for the problems $1|r_{j}|L_{\max}$ and $1|r_{j},q_{j}|C_{\max}$ was suggested in McMahon and Florian [39] and later, based on similar ideas, in Carlier [13]. A number of other special cases are solvable in polynomial time. If all jobs are of unit length, the problem is again easy to solve, see Horn [30]: indeed, while applying the ED-heuristic, no new (earlier unreleased) job can be released within the execution interval of a currently running job since job release times are integer and every job takes just a unit time of processing. As a result, no job may cause a forced delay of any more urgent job.
@@
With arbitrary release times and due dates (or tails), if all processing times are equal (but not necessarily of unit length), it is still possible to find an optimal schedule in polynomial time, see Garey et al. [25]. Even with two allowable job processing times $p$ and $2p$, for an arbitrary integer $p$, the problem $1|p_{j}\in\{p,2p\},r_{j},q_{j}|C_{\max}$ remains polynomially solvable [48]. In fact, in the algorithm from the latter reference the processing times of only some jobs (which become known during the execution of the algorithm) are restricted to $p$ or $2p$, whereas the rest of the jobs may have arbitrary processing times. The algorithm is obtained as a consequence of two enumerative algorithms for the problem $1|r_{j},q_{j}|C_{\max}$. The first algorithm is exponential. The the second algorithm is a restriction of the first one and runs in time $O(n^{2}\log n)$ under certain conditions which are to be held during its execution. Otherwise, another auxiliary procedure is to be applied to guarantee the optimality. An $O(n^{2}\log n\log P)$ implementation of that procedure for the problem $1|p_{j}\in\{P,2P\},r_{j},q_{j}|C_{\max}$ yields an algorithm with the same time complexity for this setting.
@@
$r_{i}+p_{i}\geq r_{j}+p_{j},$
$d_{i}\geq d_{j}$ holds, then the resultant problem $1|r_{j},embedded|L_{\max}$ is solvable in $O(n\log n)$ time, see Vakhania [52].
@@
For the problem $1|r_{j},p_{j}=p|L_{max}$ with constant processing times, Lazarev et al. [38] presented two polynomial algorithms. The first procedure is based on trisection search and determines an optimal solution in $O(Qn\log n)$ time, where $10^{-Q}$ represents the accuracy of the parameters of the problem. The second approach uses an auxiliary problem, where the maximum completion time is minimized under the constraint that $L_{max}$ is bounded by a particular value. Using this approach, the authors determine for the bi-criteria problem with the two criteria $C_{max}$ and $L_{max}$ and the Pareto set of optimal solutions in $O(n^{3}\log n)$ time.
With arbitrary job processing times, the problem remains (weakly) $NP$-hard already with only two allowable job release times and due dates $1|r_{j}\in\{r_{1},r_{2}\},d_{j}\in\{d_{1},d_{2}\}|L_{\max}$, see Chinos and Vakhania [14]. This is a minimal $NP$-hard special case of the problem $1|r_{j},d_{j}|L_{\max}$. The latter setting was later studied in Reynoso and Vakhania [44], where structural properties of this special case have led to a number of optimality conditions under which it can be solved optimally in $O(n\log n)$ time were established. It was shown that, for bottleneck instances where none of these conditions apply, the problem can be reduced to the SUBSET SUM problem. The authors presented computational results accomplished for 50 million randomly generated problem instances, reporting that almost all these instances were solved optimally by one of the established optimality conditions. Moreover, the heuristic algorithms that incorporate the optimality conditions established in [44] succeeded to solve the SUBSET SUM problem also optimally for almost all the 50 million instances. A very recent work considered a natural generalization of the latter problem with any fixed number of release times, see Escalona and Vakhania [17]. For this setting, the optimality conditions were derived and again, it was shown how the problem can be reduced to the SUBSET SUM. It is a challenging question to what extent the above obtained results can be generalized for any fixed number of release and delivery times, a setting with much wider practical applications.
@@
A special kind of parameterized analysis called the Variable Parameter (VP) analysis was recently suggested in [54]. Based on this approach, an implicit enumeration algorithm and a polynomial time approximation scheme for the general setting $1|r_{j},q_{j}|C_{\max}$ were proposed. The variable parameter is defined as the number of emerging jobs. First, a partial solution in time $O(n\log n)$ without these jobs is constructed. Then this partial solution is augmented to a complete optimal solution in time, which is exponential solely in the number of the variable parameter (the number of emerging jobs). The author also gives alternative time complexity expression for both, the implicit enumeration and the polynomial time approximation scheme, in which the exponential dependence is solely on the maximum processing time and the maximum delivery time of an emerging job. If one applies the fixed parameter analysis to these estimations, one gets a polynomial-time dependence. A very recent extensive experimental study showed that, in practice, the variable parameter is much less than $n$. Moreover, the ratio of the variable parameter to $n$ asymptotically converges to 0 (see Ramirez and Vakhania [43]).
The problem $1|r_{j}|L_{\max}$ is important on its own and also because of its possible applications for the solution of other more complex scheduling problems and other optimization problems. For example, in the Multiple Traveling Salesman Problem (M-TSP), $k$ salesmen, each starting from a special point 00 called depot, has to visit a number of clients and return to the depot. The objective is to minimize the total cost of all $k$ tours (see, for example [42]). In the setting with time windows M-TSP-TW, each client has to be visited within a specified time interval. In the former basic setting, it is easy to find a feasible solution to the problem. However, in the latter setting with time windows, there may exist no feasible solution for a given $k$. Hence, it is crucial to determine whether there exists a feasible solution for a given $k$. More generally, one may be interested to calculate the minimum $k$ for which there exists a feasible solution. We can use our scheduling problem for this purpose.
In the basic case, we wish to determine whether there exists a feasible solution with $k=1$. For a given instance of M-TSP-TW with $n+1$ clients, we create an instance of he problem $1|r_{j},p_{j}=0|L_{\max}$ with $n$ jobs, associating job $j$ with client $j$, for each $i=1,\dots,n$, and letting the processing time of each job to be 0. Note that this is a simplified version of the generic problem $1|r_{j}|L_{\max}$. In this version, the release and due times of each job are determined according to the time window of the corresponding client. We consider the feasibility version of the problem, where one looks for a schedule in which no job has a positive lateness. One can immediately see that, if there is no feasible solution to the scheduling problem, there is also no feasible solution for the corresponding instance of the problem M-TSP-TW with $k=1$ (i.e., TSP-TW). In general, we let $u$ be an upper bound on the total number of salesmen, and carry out binary search within the interval $[1,u]$ solving at each iteration an instance of the corresponding scheduling problem $Pk|r_{j},p_{j}=0|L_{\max}$ with $k$ parallel identical machines and with trial $k\in[1,u]$. In this way, we can find the minimum $k$ for which there exists a feasible solution to the problem M-TSP-TW.